# Characterization of coherent errors in gate layers with robustness to Pauli noise

### v2

@@ I Introduction
Demarcating noise along the line of purity conservation leads to the distinction between coherent and incoherent noise [9]. Coherent noise occurs due to systematic, reversible perturbations such as imperfect calibrations, imprecise control signals, or couplings to low-frequency external fields [10, 11, 12, 13]. Conversely, incoherent noise, e.g., bit-flip errors, is stochastic and typically arises from insufficient isolation of the system from its environment [12], causing a non-reversible loss of information. Since controllability and isolation of a system are generally opposing goals in hardware design [4, 14], distinguishing the structure of the limiting noise is critical to advance quantum computing platforms.
@@
Numerous characterization techniques have been developed to identify and quantify coherent and incoherent noise in quantum processes. Full process tomography [15, 16] and gate-set tomography [17, 18] are common techniques for obtaining a general representation of a physical operation. Although they provide a full representation, they lack in discriminating the noise processes from the actual operation. Furthermore, as they involve the reconstruction of full-system density matrices, without further assumptions, they are experimentally intractable on devices with more than a few qubits due to exponential scaling of the number of operations with the size of the system’s Hilbert space and the resource-intensive post-processing [19]. While some protocols reach scalable noise characterization using principles of randomized benchmarking [20, 21, 22], they are often limited to a few heuristic metrics that do not provide enough insights for existing noise calibration, mitigation, or correction techniques [23]. Characterization approaches that exceed this limitation while still being scalable exist for incoherent noise [24, 7, 25]. These noise reconstruction protocols [26] introduce a model and estimate its parameters by fitting it to the real process. They aim to quantify noise of a gate layer, where a restricted noise correlation length is the central structural assumption. However, those models do not portray the actual noise process completely, as the protocols require the application of probabilistic twirling schemes [27, 28, 20, 29, 30, 31] that convert coherent errors into incoherent ones. Currently, no similar approach to the noise reconstruction protocols exists that is sensitive to coherent errors.
@@
In this article, we introduce a noise model combining coherent and incoherent components. We devise a protocol to characterize the coherent part of the model as a series of unitary transformations, interpretable as qubit rotations. The estimation is independent of the introduced incoherent model components, hardware-agnostic, and scalable to larger systems, assuming a low noise-correlation length. To demonstrate our method, we apply it to characterize a gate layer on a 7-qubit superconducting circuit device provided by IBM Quantum (IBMq) [32]. This allows us to determine the most significant coherent single- and two-qubit errors and their behavior over time. Leveraging these estimation results, we mitigate the coherent errors through a gate-level approach and combine it with probabilistic error cancellation (PEC) [8, 33]. Thereby verifying that our model captures the dominating noise terms. Our work paves the way for advancements in hardware platform development, quantum error correction and mitigation schemes, and the design of noise-aware algorithms.
@@ II Model
An ideal noiseless operation on a quantum computer can be expressed by a unitary channel $\mathcal{U}_{I}(\rho)=U_{I}\rho U_{I}^{\dagger}$, with $U_{I}$ being some unitary matrix. A simple model for the noisy implementation of this operation on real hardware $\mathcal{E}_{\mathcal{P}}$ consists of the concatenation of a noise modeling Pauli channel $\mathcal{P}$ and the ideal unitary,
@@
where the Pauli error rates $p(i)$ form a probability distribution over the $n$-qubit Pauli operators $P_{i}\in\mathbb{P}^{\otimes n}=\{I,X,Y,Z\}^{\otimes n}$. Despite the substantial expressibility restrictions of the model by only considering Pauli operators, a wide range of unital incoherent noise, including dephasing and depolarization, can be described by Pauli channels. In this framework, a $n$-qubit noisy operation is described at most by $4^{n}$ Pauli error rates. To reduce this scaling from exponential to polynomial, we follow the widely used and experimentally motivated approach of only regarding locally correlated noise, i.e., considering $n$-qubit Pauli operators $P_{i}$ that act nontrivially on $l$ qubits only [24, 7]. These $l$ qubits are selected with respect to the qubits’ connectivity in the quantum computing architecture. This work focuses on a correlation length $l=2$.
@@
Pauli noise models are appealing for different physical and practical reasons. Depolarization and dephasing are often dominating the qubit systems’ interaction with the environment [34, 35]. Quantum error correction also leads to noise better approximated by Pauli channels [36, 37, 38]. Additionally, Pauli channels are interesting for theoretical work as they can be efficiently simulated classically [39]. Most importantly, however, any noise can be projected onto a Pauli channel by applying Pauli twirling [27, 28, 20, 29, 30, 31]. If $\mathcal{U}_{I}$ is a Clifford operation Pauli twirling can be achieved by stochastically introducing Pauli gates before and after the ideal operation. Enforcing the structure of the model given in Eq. (1) by twirling allows the mitigation of any unknown general noise channel with schemes tailored to Pauli noise, such as PEC. Despite the sampling and circuit overhead due to the probabilistic nature of PEC and twirling [27, 28, 24], this mitigation approach is well-suited for near-term devices because it is implementable with low-depth circuits and relies on Pauli gates [8, 24].
@@
However, in the context of characterization where the goal is to examine the physical noise sources of a system, altering the noise structure using twirling is often not desired, as it obscures the distinction between coherent and incoherent noise. To illustrate the transformative effect of Pauli twirling and motivate the extension of the Pauli noise model in Eq. (1) for characterization purposes, we perform an echo-experiment, shown in Fig. 1 (similar to Ref. [40]), on an IBMq processor. The circuits consist of preparing the state |$\ket{0+}$, an even number of $C\!X$ gates, and measuring the evolved state in the same Pauli basis as it was prepared. As two $C\!X$ gates correspond to the identity, in the noiseless case all measured expectation values are independent of the number of repetitions and equal to $1$. In the exclusive presence of Pauli noise, purely exponential decay of the expectation values, characteristic of incoherent error processes, is expected [20, 21, 22, 28, 41]. This behavior (red dotted curves) can be observed by adding Pauli twirling to the circuits. In the absence of twirling (blue dashed curves), oscillations appear, a feature that exceeds the expressibility of the Pauli noise model in Eq. (1).
@@
Figure 1: Noise examination on identity circuit. The data is obtained on ibm_lagos and plots the expectation value of the three two-qubit Pauli operators $P_{IX},P_{ZI},P_{ZX}$ against the number of noisy identities $m$ implemented by two $C\!X$ gates. In the experiment, we prepare and measure the system in the state |$\ket{0+}$ using one Hadamard gate $H$. The blue dashed curves correspond to the unchanged noise. The dotted red curves are obtained using Pauli twirling scheme that projects the noise to Pauli noise (see Eq. (1)). Each circuit is measured $4096$ times and the twirled results are averaged over $32$ different twirling gates $P$. A readout mitigation scheme [42] is applied in post-processing based on the reported readout error rates of IBM Quantum. The results for different input states and superconducting hardware platforms available on IBMq are similar.
@@
with $U_{\theta}$ being a unitary operator parameterized by a Hermitian matrix $H_{\theta}$ as ${U_{\theta}=e^{-iH_{\theta}}}$. This matrix can be expressed in the Pauli basis as $H_{\theta}=\sum_{k}\theta_{k}P_{k}$, with real coefficients $\theta_{k}$ and Pauli matrices $P_{k}$. Assuming small noise, i.e., $U_{\theta}$ is close to identity, we can rewrite the coherent noise as $U_{\theta}\approx\prod_{k}\exp\!\left(-i\theta_{k}P_{k}\right)$, which is a valid approximation up to $\mathcal{O}(\theta^{2})$. Each product term reads
@@
and represents in the Bloch sphere picture a rotation around the $k$-axis by the angle $2\,\theta_{k}$, see Fig. 2. To reach a scalable model, we consider, as for the Pauli noise model in Eq. (1), locally correlated noise operators, namely single- and two-qubit rotations among neighboring qubits according to their connectivity in the hardware. Therefore, the coherently rotated Pauli noise model defined in Eq. (2) consists of nine two-qubit rotations for each neighboring qubit pair and three single-qubit rotations for each qubit, leading to a magnitude of these rotations of $\lvert\theta\rvert=15$.
@@ III Characterization
We express the coherent noise characterization protocol in the Pauli transfer matrix (PTM) formalism [49]. As the Pauli matrices build a complete basis of the operator space, a $n$-qubit state $\rho$ can be vectorized $\left|\rho\right)\in\mathbb{R}^{4^{n}}$ in the Pauli basis with the $i$’th element $\left|\rho\right)_{i}$ being their expectation values ${\left|\rho\right)_{i}=\frac{1}{2^{n}}\mathrm{tr}\left({P_{i}\rho}\right)}$. These $4^{n}$ expectation values can be estimated from $3^{n}$ measurements of $\rho$ in different Pauli bases [50, 51, 52, 7, 53, 54]. For a single qubit, this state representation is known as the Bloch vector [55]. In this formalism, a channel $\mathcal{E}$ is represented by a matrix $T\in\mathbb{R}^{4^{n}}\times\mathbb{R}^{4^{n}}$ that acts on the vectorized states according to $\left|\mathcal{E}\!\left(\rho\right)\right)=T\left|\rho\right)$. The matrix $T$ is called the PTM, with elements $T_{ij}=\mathrm{tr}\left(P_{i}\,\mathcal{E}\!\left(P_{j}\right)\right)$. In the context of the coherently rotated Pauli noise model, the diagonal of the PTM matrix encompasses information about the incoherent Pauli noise channel mixed with the second-order terms of the coherent noise. The first-order terms of the coherent noise are contained in the off-diagonal elements without any contribution from Pauli noise.
Before addressing the characterization protocol applicable to a quantum process modeled by the coherently rotated Pauli noise model in Eq. (2), we present the case of estimating $U_{\theta}$ for a 2-qubit process described by $\mathcal{E}_{\theta}\!\left(\rho\right)=U_{\theta}\rho U_{\theta}^{\dagger}$, meaning that we disregard both the Pauli noise $\mathcal{N}$ and the ideal unitary $\mathcal{U}_{I}$ channels. The Pauli transfer matrix $T_{\theta}$ corresponding to $\mathcal{E}_{\theta}$ is a square matrix of order $16$, parameterized by $15$ single- and two-qubit rotation angles. While the first row and column are trivial and independent of $\theta$, in a first-order approximation in $\theta$ (see App. A for a second-order approximation), the other elements of this matrix are given by
@@
By experimentally estimating the vector representation of a set of states $\{\rho_{i}\}$ and their propagation $\{\mathcal{E}_{\theta}\!\left(\rho_{i}\right)\}$ we get an estimate $\theta$ by solving the minimization problem
@@
Note that compared to full-process tomography on two qubits the number of real parameters is reduced from $240$ to $15$ real parameters, and for a system with $l$-qubits from $\mathcal{O}\left(16^{l}\right)$ to $\mathcal{O}\left(4^{l}\right)$. In App. B we provide the closed-form solution of the minimization in Eq. 5 for a first-order approximation in $\theta$.
The strategy of the characterization protocol for the entire $n$-qubit system is to apply this estimation procedure to every involved two-qubit subsystem according to their connectivity in the hardware. Three points must be considered when moving from the discussed simple case to the model given in Eq. (2): first, the action of the ideal unitary $\mathcal{U}_{I}$; second, the presence of the Pauli noise $\mathcal{N}$; and lastly, the impact of two-qubit noise on qubits that are outside of the considered subsystem.
The effect of the ideal unitary channel $\mathcal{U}_{I}$ can be incorporated in post-processing after estimating the state $\left|\rho\right)$. For an arbitrary $U_{I}$, this procedure requires a full-state tomography of the system [39]. To preserve scalability, we restrict to operations, separable into multiple single- and two-qubit unitaries $U_{I}=U_{1}\otimes U_{2}\otimes...\otimes U_{l}$. This requirement is fulfilled, e.g., in the characterization of so-called gate layers, i.e., $n$-qubit operations consisting of single- and two-qubit gates that can be executed simultaneously in experiments.
@@
Lastly, when focusing exclusively on two-qubit subsystems, one has to address the correlated multi-qubit noise between a qubit in the regarded subsystem and the external ones. To handle this problem in the estimation of $\left|\rho\right)$ and $\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)$, we prepare the surrounding qubits in different basis states for every single experiment run. This technique isolates the two desired qubits and randomizes the effect of the noise introduced outside the regarded subsystem, similar to twirling schemes.
The protocol for the estimation of coherent noise for a two-qubit system is as follows (for more than 2-qubits see App. D):
@@
    Choose a positive integers $j$.
@@
    Initialize the $n$-qubit system in |$\ket{0}^{\otimes n}$.
@@
    Prepare the system in the eigenstate of a random $n$-qubit Pauli operator $P$ by applying to each qubit $H$, $HX$, $SH$, $SHX$, $X$ or $\mathbb{I}$, obtaining a state |$\ket{\rho}$, with $P\ket{\rho}=\pm\ket{\rho}$.
@@
    Aggregate all pairs $\left\{\left|\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)\right),\left|\mathcal{U}^{\dagger}_{I}\,\mathcal{E}_{\theta}^{j+1}\!\left(\rho_{j}\right)\right)\right\}$ for all prepared states and repetitions $j$ and perform the minimization of Eq. (5).
@@
Figure 3: Characterization of the coherent noise introduced by executing a gate layer on the $7$-qubit IBMq Processor ibm_lagos [32]. The gate layer is drawn on top of the connectivity tree (qubit connectivity in the quantum computing architecture), where each big circle marks a different qubit. The gate layer consists of two $C\!X$ gates between qubits $1$ and $2$, and qubits $6$ and $7$, a Hadamard ($H$) gate on qubit 3, a $\sqrt{X}$ ($S\!X$) gate on qubit $4$ and a $X$ gate on qubit $5$. The three pieces inside the qubit representing circles represent the single-qubit coherent errors, and the nine small circles between two qubits display the two-qubit coherent cross-talk errors. All errors are measured in radians and the absolute value is plotted.
@@
We now use the presented protocol to characterize the coherent noise induced by an arbitrary gate layer executed on the IBMq Falcon Processor ibm_lagos. The result is summarized in Fig. 3, where we show the absolute values of the angles of the characterized single- and two-qubit coherent errors for an arbitrary gate layer. The layer is composed of two $C\!X$ gates and three single-qubit gates, namely $H$, $\sqrt{X}$, and $X$. Considering the connectivity of the $7$ qubits in the hardware and our assumption that the noise is only locally correlated, we prepare $216$ initial states, all being product states of |$\ket{0}$, |$\ket{1}$, |+=(|$\ket{+}=(\ket{0}+\ket{1})/\sqrt{2}$, |=(|$\ket{-}=(\ket{0}-\ket{1})/\sqrt{2}$, |$\ket{i}=(\ket{0}+i\ket{1})/\sqrt{2}$ and |$\ket{-i}=(\ket{0}-i\ket{1})/\sqrt{2}$. Each two-qubit subsystem is prepared in $36$ states, and for each of those states, the environment is randomized with $6$ additional states. We execute 00 to $3$ repetitions of the layer, amounting to $4$ circuits per preparation. Finally, bypassing the effect of the unitary gate layer $\mathcal{U}_{I}$ in the post-processing necessitates performing state tomography on up to three-qubit subsystems, leading to $3^{3}$ Pauli measurements per circuit. Overall, we run $328216\cdot 4\cdot 27=23^{\prime}$ 328 different circuits with $128$ shots each, leading to a total running time in the system of about $30$ minutes. A more detailed description of the procedure is given in App. D.
@@
Figure 4: Noise mitigation according to the coherently rotated Pauli noise model in Eq. (2) on ibm_lagos. The plot shows the estimated expectation values of all 9 two-qubit Pauli operators for the initial state $\ket{++}$ after $m$ repetitions of a noisy identity (a) without mitigation, (b) with coherent error mitigation based on data of (a), and (c) with probabilistic error cancellation based on data of (b) and coherent error mitigation. $\ket{++}$ is an eigenstate of the operators $P_{I\!X}$, $P_{X\!I}$, and $P_{X\!X}$, in the noiseless case, their expectation values are $1$. The expectation values of the other 12 Pauli operators (red) are 00. The circuit diagrams contain the preparation of each qubit in one of the states |${\ket{0},\ket{+},\ket{i}}$, by applying Hadamard ($H$) and/or phase ($S$) gates. The evolved states are measured in $9$ bases to estimate all $16$ Pauli expectation values. In (c), specific Pauli operators ($P$) are inserted between the circuit elements according to the PEC method.
@@
The used IBMq platforms report readout errors on the order of 1%, caused by relaxation, imperfect coupling to the readout resonator, and signal amplification errors [56]. To counter the influence of these errors, we apply a post-processing readout mitigation scheme in all executed experiments [42]. While the introduced protocol is not resistant to general state preparation and measurement errors, probabilistic errors that can be modeled by a Pauli channel do not disturb the characterization for the same reasons that the coherent noise characterization is robust against Pauli noise. This is described in detail in App. C. Furthermore, systematic, coherent, preparation errors also don’t affect the protocol, as can be seen in Eq. 5 where the estimation does not depend on information about the intended prepared state.
@@
From Fig. 3, we conclude that the most significant two-qubit coherent errors occur between qubits $1$ and $2$ and qubits $6$ and $7$. Those are associated with the Pauli $P_{Y\!Z}$ error. This result is expected, as those are the pairs of qubits on which the $C\!X$ gates are acting. This Pauli error has already been noted in the analysis of the echoed cross-resonance gate used to realize the $C\!X$ gate [57]. The predominant coherent errors are the single-qubit ones. We note no strong bias towards a specific single-qubit rotation error, and those errors could be related to the hardware implementation of each single-qubit gate.

### v3

@@ I Introduction
Demarcating noise along the lines of purity conservation leads to the distinction between coherent and incoherent noise [9]. Coherent noise occurs due to systematic, reversible perturbations such as imperfect calibrations, imprecise control signals, or couplings to low-frequency external fields [10, 11, 12, 13]. Conversely, incoherent noise (e.g., bit-flip errors) is stochastic and typically arises from insufficient isolation of the system from its environment [12], causing a nonreversible loss of information. Since controllability and isolation of a system are generally opposing goals in hardware design [4, 14], distinguishing the structure of the limiting noise is critical to the advancement of quantum computing platforms.
@@
Numerous characterization techniques have been developed to identify and quantify coherent and incoherent noise in quantum processes. Full-process tomography [15, 16] and gate-set tomography [17, 18] are common techniques for obtaining a general representation of a physical operation. Although they provide a full representation, they lack discrimination of the noise processes from the actual operation. Furthermore, as they involve the reconstruction of full-system density matrices, without further assumptions, they are experimentally intractable on devices with more than a few qubits due to exponential scaling of the number of operations with the size of the system’s Hilbert space and the resource-intensive postprocessing [19]. While some protocols achieve scalable noise characterization using principles of randomized benchmarking [20, 21, 22], they are often limited to a few heuristic metrics that do not provide enough insights for existing noise calibration, mitigation, or correction techniques [23]. Characterization approaches that exceed this limitation while still being scalable exist for incoherent noise [24, 7, 25]. These noise reconstruction protocols [26] introduce a model and estimate its parameters by fitting it to the real process. They aim to quantify noise of a gate layer, where a restricted noise correlation length is the central structural assumption. However, those models do not portray the actual noise process completely, as the protocols require the application of probabilistic twirling schemes [27, 28, 20, 29, 30, 31] that convert coherent errors into incoherent ones. Currently, no similar approach to noise reconstruction protocols exists that is sensitive to coherent errors.
@@
In this paper, we introduce a noise model combining coherent and incoherent components. We devise a protocol to characterize the coherent part of the model as a series of unitary transformations, interpretable as qubit rotations. The estimation is independent of the incoherent model components introduced, hardware-agnostic, and scalable to larger systems, assuming a low noise-correlation length. To demonstrate our method, we apply it to characterize a gate layer on a seven-qubit superconducting circuit device provided by IBM Quantum (IBMq) [32]. This allows us to determine the most significant coherent single- and two-qubit errors and their behavior over time. Leveraging these estimation results, we mitigate the coherent errors through a gate-level approach and combine it with probabilistic error cancelation (PEC) [8, 33] in order to verify that our model captures the dominating noise terms. Our work paves the way for advancements in hardware platform development, quantum error correction and mitigation schemes, and the design of noise-aware algorithms.
@@ II Model
An ideal noiseless operation on a quantum computer can be expressed by a unitary channel $\mathcal{U}_{I}(\rho)=U_{I}\rho U_{I}^{\dagger}$, for some unitary matrix $U_{I}$. A simple model for the noisy implementation of this operation on real hardware $\mathcal{E}_{\mathcal{P}}$ consists of the concatenation of a noise modeling Pauli channel $\mathcal{P}$ and the ideal unitary,
@@
where the Pauli error rates $p(i)$ form a probability distribution over the $n$-qubit Pauli operators $P_{i}\in\mathbb{P}^{\otimes n}=\{I,X,Y,Z\}^{\otimes n}$. Despite the substantial expressibility restrictions of the model by only considering Pauli operators, a wide range of unital incoherent noise, including dephasing and depolarization, can be described by Pauli channels. In this framework, an $n$-qubit noisy operation is described at most by $4^{n}$ Pauli error rates. To reduce this scaling from exponential to polynomial, we follow the widely used and experimentally motivated approach of only regarding locally correlated noise, that is, considering $n$-qubit Pauli operators $P_{i}$ that act nontrivially on $l$ qubits only [24, 7]. These $l$ qubits are selected with respect to the qubits’ connectivity in the quantum computing architecture. This work focuses on a correlation length $l=2$.
@@
Pauli noise models are appealing for different physical and practical reasons. Depolarization and dephasing often dominate the qubit systems’ interaction with the environment [34, 35]. Quantum error correction also leads to noise better approximated by Pauli channels [36, 37, 38]. Additionally, Pauli channels are interesting for theoretical work as they can be efficiently simulated classically [39]. Most importantly, however, any noise can be projected onto a Pauli channel by applying Pauli twirling [27, 28, 20, 29, 30, 31]. If $\mathcal{U}_{I}$ is a Clifford operation Pauli twirling can be achieved by stochastically introducing Pauli gates before and after the ideal operation. Enforcing the structure of the model given in Eq. (1) by twirling allows the mitigation of any unknown general noise channel with schemes tailored to Pauli noise, such as PEC. Despite the sampling and circuit overhead due to the probabilistic nature of PEC and twirling [27, 28, 24], this mitigation approach is well suited for near-term devices because it is implementable with low-depth circuits and relies on Pauli gates [8, 24].
@@
However, in the context of characterization where the goal is to examine the physical noise sources of a system, altering the noise structure using twirling is often not desired, as it obscures the distinction between coherent and incoherent noise. To illustrate the transformative effect of Pauli twirling and motivate the extension of the Pauli noise model in Eq. (1) for characterization purposes, we perform an echo experiment, shown in Fig. 1 (similar to Ref. [40]), on an IBMq processor. The circuits consist of preparing the state |$\ket{0+}$, an even number of $C\!X$ gates, and measuring the evolved state in the same Pauli basis as that in which it was prepared. As two $C\!X$ gates correspond to the identity, in the noiseless case all measured expectation values are independent of the number of repetitions and equal to $1$. In the exclusive presence of Pauli noise, purely exponential decay of the expectation values, characteristic of incoherent error processes, is expected [20, 21, 22, 28, 41]. This behavior (red dotted curves) can be observed by adding Pauli twirling to the circuits. In the absence of twirling (blue dashed curves), oscillations appear, a feature that exceeds the expressibility of the Pauli noise model in Eq. (1).
@@
Figure 1: Noise examination on identity circuit. The data are obtained on ibm_lagos, and the expectation value of the three two-qubit Pauli operators $P_{IX},P_{ZI},P_{ZX}$ is plotted against the number of noisy identities $m$ implemented by two $C\!X$ gates. In the demonstration, we prepare and measure the system in the state |$\ket{0+}$ using one Hadamard gate $H$. The blue dashed curves correspond to the unchanged noise. The dotted red curves are obtained using a Pauli twirling scheme that projects the noise to Pauli noise (see Eq. (1)). Each circuit is measured $4096$ times and the twirled results are averaged over $32$ different twirling gates $P$. A readout mitigation scheme [42] is applied in postprocessing based on the reported readout error rates of IBM Quantum. The results for different input states and superconducting hardware platforms available on IBMq are similar.
@@
where $U_{\theta}$ is a unitary operator parameterized by a Hermitian matrix $H_{\theta}$ as ${U_{\theta}=e^{-iH_{\theta}}}$. This matrix can be expressed in the Pauli basis as $H_{\theta}=\sum_{k}\theta_{k}P_{k}$, with real coefficients $\theta_{k}$ and Pauli matrices $P_{k}$. Assuming small noise, i.e., $U_{\theta}$ is close to identity, we can rewrite the coherent noise as $U_{\theta}\approx\prod_{k}\exp\!\left(-i\theta_{k}P_{k}\right)$, which is a valid approximation up to $\mathcal{O}(\theta^{2})$. Each product term reads
@@
and represents in the Bloch sphere picture a rotation around the $k$ axis by the angle $2\,\theta_{k}$, see Fig. 2. To obtain a scalable model, we consider, as for the Pauli noise model in Eq. (1), locally correlated noise operators, namely single- and two-qubit rotations among neighboring qubits according to their connectivity in the hardware. Therefore, the coherently rotated Pauli noise model defined in Eq. (2) consists of nine two-qubit rotations for each neighboring qubit pair and three single-qubit rotations for each qubit, leading to a magnitude of these rotations of $\lvert\theta\rvert=15$.
@@ III Characterization
We express the coherent noise characterization protocol in the Pauli transfer matrix (PTM) formalism [49]. As the Pauli matrices constitute a complete basis of the operator space, a $n$-qubit state $\rho$ can be vectorized $\left|\rho\right)\in\mathbb{R}^{4^{n}}$ in the Pauli basis, with the $i$th element $\left|\rho\right)_{i}$ being their expectation values ${\left|\rho\right)_{i}=\frac{1}{2^{n}}\mathrm{tr}\left({P_{i}\rho}\right)}$. These $4^{n}$ expectation values can be estimated from $3^{n}$ measurements of $\rho$ in different Pauli bases [50, 51, 52, 7, 53, 54]. For a single qubit, this state representation is known as the Bloch vector [55]. In this formalism, a channel $\mathcal{E}$ is represented by a matrix $T\in\mathbb{R}^{4^{n}}\times\mathbb{R}^{4^{n}}$ that acts on the vectorized states according to $\left|\mathcal{E}\!\left(\rho\right)\right)=T\left|\rho\right)$. The matrix $T$ is called the PTM, with elements $T_{ij}=\mathrm{tr}\left(P_{i}\,\mathcal{E}\!\left(P_{j}\right)\right)$. In the context of the coherently rotated Pauli noise model, the diagonal of the PTM matrix encompasses information about the incoherent Pauli noise channel mixed with the second-order terms of the coherent noise. The first-order terms of the coherent noise are contained in the off-diagonal elements without any contribution from Pauli noise.
Before addressing the characterization protocol applicable to a quantum process modeled by the coherently rotated Pauli noise model in Eq. (2), we present the estimation if $U_{\theta}$ for a two-qubit process described by $\mathcal{E}_{\theta}\!\left(\rho\right)=U_{\theta}\rho U_{\theta}^{\dagger}$, meaning that we disregard both the Pauli noise $\mathcal{N}$ and the ideal unitary $\mathcal{U}_{I}$ channels. The Pauli transfer matrix $T_{\theta}$ corresponding to $\mathcal{E}_{\theta}$ is a square matrix of order $16$, parameterized by $15$ single- and two-qubit rotation angles. While the first row and column are trivial and independent of $\theta$, in a first-order approximation in $\theta$ (see App. A for a second-order approximation), the other elements of this matrix are given by
@@
By estimating the vector representation of a set of states $\{\rho_{i}\}$ and their propagation $\{\mathcal{E}_{\theta}\!\left(\rho_{i}\right)\}$ we get an estimate of $\theta$ by solving the minimization problem
@@
Note that, compared to full-process tomography on two qubits, the number of real parameters is reduced from $240$ to $15$, and for a system with $l$ qubits from $\mathcal{O}\left(16^{l}\right)$ to $\mathcal{O}\left(4^{l}\right)$. In App. B we provide the closed-form solution of the minimization in Eq. 5 for a first-order approximation in $\theta$.
The strategy of the characterization protocol for the entire $n$-qubit system is to apply this estimation procedure to every two-qubit subsystem involved according to their connectivity in the hardware. Three points must be considered when moving from the simple case discussed to the model given in Eq. (2): first, the action of the ideal unitary $\mathcal{U}_{I}$; second, the presence of the Pauli noise $\mathcal{N}$; and lastly, the impact of two-qubit noise on qubits that are outside of the considered subsystem.
The effect of the ideal unitary channel $\mathcal{U}_{I}$ can be incorporated in postprocessing after estimating the state $\left|\rho\right)$. For an arbitrary $U_{I}$, this procedure requires a full-state tomography of the system [39]. To preserve scalability, we restrict to operations, separable into multiple single- and two-qubit unitaries $U_{I}=U_{1}\otimes U_{2}\otimes...\otimes U_{l}$. This requirement is fulfilled, for example, in the characterization of so-called gate layers, that is, $n$-qubit operations consisting of single- and two-qubit gates that can be executed simultaneously.
@@
Lastly, when focusing exclusively on two-qubit subsystems, one has to address the correlated multiqubit noise between a qubit in the subsystem of interest and the external ones. To handle this problem in the estimation of $\left|\rho\right)$ and $\left|\mathcal{E}_{\theta}\!\left(\rho\right)\right)$, we prepare the surrounding qubits in different basis states for every single run. This technique isolates the two desired qubits and randomizes the effect of the noise introduced outside the subsystem of interest, similar to twirling schemes.
The protocol for the estimation of coherent noise for a two-qubit system is as follows (for more than two qubits see App. D):
@@
    Choose a positive integer $j$.
@@
    Initialize the two-qubit system in |$\ket{0}^{\otimes 2}$.
@@
    Prepare the system in the eigenstate of a random two-qubit Pauli operator $P$ by applying to each qubit $H$, $HX$, $SH$, $SHX$, $X$ or $\mathbb{I}$, obtaining a state |$\ket{\rho}$, with $P\ket{\rho}=\pm\ket{\rho}$.
@@
    Aggregate all pairs $\left\{\left|\mathcal{E}_{\theta}^{j}\!\left(\rho_{j}\right)\right),\left|\mathcal{U}^{\dagger}_{I}\,\mathcal{E}_{\theta}^{j+1}\!\left(\rho_{j}\right)\right)\right\}$ for all prepared states and repetitions $j$ and carry out the minimization of Eq. (5).
@@
Figure 3: Characterization of the coherent noise introduced by executing a gate layer on the seven-qubit IBMq Falcon processor ibm_lagos [32] (specifications in App. F). The gate layer is drawn on top of the connectivity tree (qubit connectivity in the quantum computing architecture), where each big circle marks a different qubit. The gate layer consists of two $C\!X$ gates between qubits $1$ and $2$, and qubits $6$ and $7$, a Hadamard ($H$) gate on qubit 3, a $\sqrt{X}$ ($S\!X$) gate on qubit $4$, and an $X$ gate on qubit $5$. The three pieces inside the qubit representing circles represent the single-qubit coherent errors, and the nine small circles between two qubits display the two-qubit coherent cross-talk errors. All errors are measured in radians and the absolute value is plotted.
@@
We now use the above protocol to characterize the coherent noise induced by an arbitrary gate layer executed on the IBMq Falcon processor ibm_lagos. The system specifications are listed in App. F. The result is summarized in Fig. 3, where we show the absolute values of the angles of the characterized single- and two-qubit coherent errors for an arbitrary gate layer. The layer is composed of two $C\!X$ gates and three single-qubit gates, namely $H$, $\sqrt{X}$, and $X$. Considering the connectivity of the seven qubits in the hardware and our assumption that the noise is only locally correlated, we prepare $216$ initial states, all being product states of |$\ket{0}$, |$\ket{1}$, |+=(|$\ket{+}=(\ket{0}+\ket{1})/\sqrt{2}$, |=(|$\ket{-}=(\ket{0}-\ket{1})/\sqrt{2}$, |$\ket{i}=(\ket{0}+i\ket{1})/\sqrt{2}$ and |$\ket{-i}=(\ket{0}-i\ket{1})/\sqrt{2}$. Each two-qubit subsystem is prepared in $36$ states, and for each of those states, the environment is randomized with six additional states. We execute 00 to $3$ repetitions of the layer, amounting to four circuits per preparation. Finally, bypassing the effect of the unitary gate layer $\mathcal{U}_{I}$ in the postprocessing necessitates performing state tomography on up to three-qubit subsystems, leading to $3^{3}$ Pauli measurements per circuit. Overall, we run $216\cdot 4\cdot 27=23328$ different circuits with $128$ shots each, leading to a total running time in the system of about $30$ minutes. A more detailed description of the procedure is given in App. D.
@@
Figure 4: Noise mitigation according to the coherently rotated Pauli noise model in Eq. (2) on ibm_lagos. The plot shows the estimated expectation values of all 15 two-qubit Pauli operators (excluding $\mathbb{I}\otimes\mathbb{I}$) for the initial state $\ket{++}$ after $m$ repetitions of a noisy identity (a) without mitigation, (b) with coherent error mitigation based on data of (a), and (c) with probabilistic error cancelation based on data of (b) and coherent error mitigation. $\ket{++}$ is an eigenstate of the operators $P_{I\!X}$, $P_{X\!I}$, and $P_{X\!X}$, in the noiseless case, and their expectation values are $1$. The expectation values of the other 12 Pauli operators (red) are 00. The circuit diagrams contain the preparation of each qubit in one of the states |${\ket{0},\ket{+},\ket{i}}$, by applying Hadamard ($H$) and/or phase ($S$) gates. The evolved states are measured in nine bases to estimate the $15$ Pauli expectation values. In (c), specific Pauli operators ($P$) are inserted between the circuit elements according to the PEC method.
@@
The IBMq platforms used report readout errors of the order of 1%, caused by relaxation, imperfect coupling to the readout resonator, and signal amplification errors [56]. To counter the influence of these errors, we apply a postprocessing readout mitigation scheme in all demonstrations [42]. While our protocol is not resistant to general state preparation and measurement errors, probabilistic errors that can be modeled by a Pauli channel do not disturb the characterization for the same reasons that the coherent noise characterization is robust against Pauli noise. This is described in detail in App. C. Furthermore, systematic, coherent, preparation errors also do not affect the protocol, as can be seen in Eq. 5 where the estimation does not depend on information about the intended prepared state.
@@
From Fig. 3, we conclude that the most significant two-qubit coherent errors occur between qubits $1$ and $2$ and between qubits $6$ and $7$. Those are associated with the Pauli $P_{Y\!Z}$ error. This result is expected, as those are the pairs of qubits on which the $C\!X$ gates are acting. This Pauli error has already been noted in the analysis of the echoed cross-resonance gate used to realize the $C\!X$ gate [57]. The predominant coherent errors are the single-qubit ones. We note no strong bias towards a specific single-qubit rotation error, and those errors could be related to the hardware implementation of each single-qubit gate.